OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of f[g(−4)]? −12 3 −13 8

ganeshie8 (ganeshie8):

first find \(g(-4)\)

ganeshie8 (ganeshie8):

\(g(x) = x+5\) \(g(-4)=?\)

OpenStudy (anonymous):

x-20?

ganeshie8 (ganeshie8):

Nope, you need to replace "x" by "-4"

OpenStudy (anonymous):

I am confused.

ganeshie8 (ganeshie8):

Easy, \(g(x) = x+5\) \(g(-4)=-4+5 = 1\)

ganeshie8 (ganeshie8):

\(f(g(-4)) = f(1) = ?\)

ganeshie8 (ganeshie8):

find the value of \(f(1)\) and you're done!

OpenStudy (anonymous):

Is it D?????? 8?????????????

ganeshie8 (ganeshie8):

nope, don't guess haha

OpenStudy (anonymous):

._.

ganeshie8 (ganeshie8):

\(f(x)=x^2-13\) \(f(1)=?\)

ganeshie8 (ganeshie8):

replace \(x\) by \(1\) on right hand side expression

OpenStudy (anonymous):

So -12?????

ganeshie8 (ganeshie8):

Yes!

OpenStudy (anonymous):

THANK YOU!!!!!!

ganeshie8 (ganeshie8):

np

OpenStudy (anonymous):

@ganeshie8 can you check 1 for me?

OpenStudy (anonymous):

If (9x − 4)(9x + 4) = ax2 − b, what is the value of a? The value of a is _____________. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

Is it 81??

ganeshie8 (ganeshie8):

Correct!

OpenStudy (anonymous):

Oh okay thank you so much!

ganeshie8 (ganeshie8):

yw